OpenStudy (shelld93):

Solve for t, include all working 3t - 6 = -2t + 6

OpenStudy (anonymous):

add 2t+6 to both sides of the equation

OpenStudy (shelld93):

3t - 6 = -2t + 6 3t - 6 + 2t + 6 = -2t + 6 + 2t + 6 3t + 2t - 6 + 6 = 6 + 6 5t - 6 + 6 = 12 5t= 12 5t/5 = 12/5 t = 12/5 correct ?

OpenStudy (anonymous):

Yes it is

OpenStudy (shelld93):

okay so to prove this id substitute 12/5 for t in the equation and each side should be the same number when solved

OpenStudy (skullpatrol):

Yes

OpenStudy (anonymous):

yeah and depending on what form they want the answer it is 12/5 or 2.4

OpenStudy (shelld93):

@skullpatrol Sorry I had to close the other one, I couldnt keep up with other people jumping in and deleting replies. Im no good a maths what so ever so its hard to me to follow.

OpenStudy (shelld93):

i tried with 2.4 before and i didnt get the same answer but i did with 12/5 did i just make a mistake ?

OpenStudy (anonymous):

so the answer is t=12/5

OpenStudy (shelld93):

thank you :)

OpenStudy (skullpatrol):

Thanks for asking :)

OpenStudy (shelld93):

thank you both :)

OpenStudy (anonymous):

you're welcome :)